# IFT-UAM/CSIC-24-92

### v1

@@ Appendix A The case of a single large dimension and Yukawa couplings
One thus obtains neutrino gonions at $10$ MeV and $M_{s}=10^{8}$ GeV. Its Yukawa will be $Y\simeq g_{\nu}^{2}\simeq 10^{-20}$, and the Dirac neutrino masses would be too small by a factor of $10^{8}$. Additionally, the rest of the SM Yukawas would be much suppressed, as argued above.
General case interpolating between one and two large dimensions
More generally one can consider limits of the form: [^7] For $\lambda>1$ we get into a non-perturbative regime where $m_{\rm D0}\simeq M_{\rm P}/t^{1/2}\simeq M_{\rm P}/u^{\lambda/2}$ becomes lighter than the string scale, the M-theory dimension emerges, and we no longer have a single large extra-dimension. The case $\lambda=1$ is borderline between perturbative and non-perturbative limits.:
$s,u^{(2)}\ \longrightarrow\ u\ ;\ t\ \longrightarrow\ u^{\lambda}\ ,\ \lambda\in[0,1]$ (A.10)
Note that for the 10d dilaton $e^{-\phi}\rightarrow u^{(1-\lambda)/2}$ so that $e^{\phi_{4}}\rightarrow 1/u^{1/2}$ and $R_{3}\rightarrow u^{(1+\lambda)/2}$. Then one can obtain interpolating solutions with:
$m_{\rm KK}\,\simeq\,\frac{M_{\rm P}}{u^{1+\lambda/2}}\,,\qquad m_{\rm w}\,\simeq\,\frac{M_{\rm P}}{u^{1-\lambda/2}}\,.$ (A.11)
Duality exchanges $\lambda\leftrightarrow-\lambda$. Gauge couplings and gonion masses keep the same expression:
$g^{2}\,\simeq\,\frac{1}{u}\,,\qquad m_{\rm gon}\,\simeq\,\frac{M_{\rm P}}{u}\,\simeq\,g^{2}M_{\rm P}\,,$ (A.12)
and $M_{s}=gM_{\rm P}$. One can also write
$m_{\rm KK}\,\simeq\,g^{2+\lambda}M_{\rm P}\,\qquad m_{\rm w}\,\simeq\,g^{2-\lambda}M_{\rm P}\,.$ (A.13)
The Yukawa coupling has the behaviour
$Y\,\simeq\,\frac{1}{t^{\lambda/2}}\frac{1}{u^{1/2}}\,\simeq\,\frac{1}{u^{\frac{1+\lambda}{2}}}\,\simeq\,g^{1+\lambda}\,,$ (A.14)
so that the gauge coupling is related to the Yukawa coupling by
$g\,\simeq\,Y^{\frac{1}{1+\lambda}}\,.$ (A.15)
Notice the mass inequalities:
$m_{\rm KK}\,\leq\,gYM_{\rm P}\,\leq\,YM_{\rm P}\,,\quad m_{\rm gon}\,=\,g^{2}M_{\rm P}\,\leq\,g^{1+\lambda}M_{\rm P}\,\leq\,YM_{\rm P}\,.$ (A.16)
Hence, indeed as the Yukawa coupling goes to zero, both KK and gonion towers appear, although at different rates, unless $\lambda=0$. Moreover one always has $m\leq YM_{\rm P}$, so that Yukawa interaction is always stronger than the gravitational interaction.
If we apply this structure to neutrino masses in which $Y\simeq 10^{-12}$ one finds
[…]
which varies in the range $10^{-12}\leq g\leq 10^{-6}$. So there is a family of directions able to reproduce the neutrino data, with values of the string scale in the range $10^{6}\,{\rm GeV}\leq M_{s}\leq 10^{12}$ GeV, and neutrino gonions in the range $10^{3}$ eV $\leq 10^{6}$ GeV respectively. However, as we said, the possibilities with $\lambda>0$ yield too small Yukawas for the rest of the SM fermions and hence are not viable.
Generalisation to CY geometries
[…]
Our premise will be the following: in order to avoid having more than one KK or winding tower coming from the Calabi–Yau geometry, we send to infinity a Kähler modulus $t$ at the same rate as the leading complex structure modulus $u$, both of them growing to infinity. Then, to avoid an overall suppression of Yukawa couplings, we only consider trajectories in Kähler moduli space such that VolX remains constant. We are then led to a very similar setup to the one considered in [84], except that we have an orientifold projection. One effect of this projection is to remove those directions in Kähler moduli space that correspond to harmonic two-forms that are even under the action of ${\cal R}$, although our analysis will be insensitive to this fact.
The classification of infinite-distance limits in Kähler moduli space of a CY performed in [84] is based on a theorem stating that, to keep VolX constant, the CY $X_{6}$ should display a fibration structure. This fibration is either of form:
A)   a ${\bf T}^{2}_{f}$ fibered over a four-dimensional base $B_{4,b}$  ,
B)   a $B_{4,f}={\bf K3}$ or ${\bf T}^{4}$ fibered over a $\mathbb{P}^{1}_{b}$  .
[…]
In case A) one grows a Kähler modulus $t$ of the base $B_{4,b}$. Then it may happen that ${\rm Vol}(B_{4})$ grows either like $t$ or $t^{2}$. In the first case we are secretly in case B) [84], so let us focus on the second case. To keep the volume constant, we have that ${\rm Area}({\bf T}^{2})\sim t^{-2}$. From here, one has at least one KK-like tower made up of D2-branes wrapping the ${\bf T}^{2}$ multiple times, [^8] Unlike in the plain Calabi–Yau setup, due to the orientifold projection the D2-brane tower is now made of non-BPS states. However just like for a tower of D0-branes, we still expect an infinite tower of states to be present (see footnote 4). whose mass scale is given by
$m_{\rm D2}\sim\frac{M_{s}}{t^{2}}\sim\frac{e^{\phi_{4}}}{u^{2}}M_{\rm P}\sim\frac{M_{\rm P}}{u^{2+q/4}}\,,$ (A.18)
where we have assumed that $e^{\phi_{4}}\sim u^{-q/4}$, $q=1,2,3,4$ along the limit. So we get an additional large extra dimension competing with the initial one. Additionally, we could have further KK-like towers coming from the winding modes of the shrinking ${\bf T}^{2}$.
Turning to case B), we now have a $\mathbbm{P}^{1}$ base that grows its area like $t$ and a fibre $B_{4,f}$ that shrinks its volume like $1/t$. We moreover assume that the limit $t\sim u\to\infty$ takes the $\mathbbm{P}^{1}$ to a spheroid whose area grows to infinity while its eccentricity approaches 1, so that one effectively recovers a single light KK scale [85]. If things work similarly to the toroidal case, one expects such a KK tower to scale as in (A.6).
Now, in the case that $B_{4,f}={\bf K3}$, there should be an obstruction to shrink the quantum volume of $B_{4,f}$ to zero. This was argued in [84] for type IIA CY compactifications without orientifolds, but one would expect the same obstruction to arise with the orientifold projection as well. As a result, there is no way to keep the quantum volume of $X_{6}$ constant, which is what should appear in the denominator of (A.4) in this regime. For $B_{4,f}={\bf T}^{4}$ one can shrink the fibre to zero size, and the leading D-brane tower made up from D4-branes multi-wrapped on $B_{4,f}$, with mass scale
$m_{\rm D4}\sim\frac{{\rm Vol}(B_{4,f})}{{\rm Vol}_{X}^{1/2}}M_{\rm P}\sim\frac{M_{\rm P}}{u}\,.$ (A.19)
Therefore, we will find at least two KK towers below the species scale, unless the latter scales like $\Lambda_{\rm QG}\sim u^{-1}M_{\rm P}$. This seems however unlikely in this particular setup, since the species scale is at most the string scale, and one expects a behaviour similar to toroidal orbifold setups, where $M_{s}\sim u^{-1/2}M_{\rm P}$.
We therefore conclude that limits of large complex structure and Kähler moduli at fixed total volume face serious difficulties in reproducing only one large extra dimension below the species scale.
Difficulties with more than two large dimensions
@@
One may also think of considering models with more than two large dimensions. One can see however that if we were able to build a model leading to tiny neutrino Yukawa couplings with more than two large dimensions, the string scale would be lowered much below the LHC threshold. We will just give a heuristic idea which goes as follows. The species scale $\Lambda_{\rm QG}$ (which is our setup is the string scale) generated by KK-towers of characteristic mass $m_{t}$ is computed in Planck units as:
@@
If we now set the value of the Yukawa to $Y_{\nu}\simeq 7\times 10^{-13}$, for $p=4,6$ one obtains a species scales of order $100$, 1,\ 1, 1 GeV respectively, which are obviously ruled out experimentally. We have not considered an odd number of dimensions because we have already seen that this requires giving large values to the Kähler moduli, which is problematic as it leads to either additional towers or too small Yukawa couplings for all fermions.

### v2

@@ A.2 One large dimension and the ‘Dark Dimension’ scenario
where we have taken conservatively $g_{\rm SM}\sim 1$ (if SUSY is not assumed gonions in these sectors could be as light as $\simeq 10^{-10}M_{s}$, which would be ruled out). With $M_{s}\simeq 10^{8}$ GeV, this means that up to 4 new extra large dimensions should develop at a scale $\lesssim 1$ TeV or below, which is already ruled out experimentally. Furthermore the SM gauge couplings would be too small due to the contribution to the running of gauge couplings from the towers of SM gonions.
So as a second summary: a single large dimension along a direction which features light $\nu_{R}$ gonions is consistent. However if the gonions are associated to the SM particles, it is not viable
Some difficulties with more than two large dimensions
@@
One may also think of considering models with more than two large dimensions. One can see however that if we were able to build a model leading to tiny neutrino Yukawa couplings with more than two large dimensions, the string scale would be lowered much below the LHC threshold. We will just give a heuristic idea which goes as follows. The species scale $\Lambda_{\rm QG}$ (which is our setup is the string scale) generated by KK-towers of characteristic mass $m_{t}$ is computed in Planck units as (see e.g. [70])
@@
If we now set the value of the Yukawa to $Y_{\nu}\simeq 7\times 10^{-13}$, for $p=4,6$ one obtains a species scales of order $100$, 1,\ 1, 1 GeV respectively, which are obviously ruled out experimentally.